OpenStudy (anonymous):

The particular solution of the differential equation dy /dt = 2y for which y(0) = 60 is a: y = 60e^2t b: y = 60 e^0.5t c: y = 59 + e^t d: y = 30e^t

OpenStudy (anonymous):

Firstly bring \(2y\) on left hand side..

OpenStudy (anonymous):

\[\frac{dy}{dt} - 2y = 0\]

OpenStudy (anonymous):

ok what happens next? :)

OpenStudy (nirmalnema):

a: y = 60e^2t

OpenStudy (anonymous):

It is basically a Linear Equation : It is precisely of the form in general: \[\frac{dy}{dt} + Py = Q\] P and Q are any function of \(t\).

OpenStudy (anonymous):

@nirmalnema you are going against Code of Conduct of this site..

OpenStudy (anonymous):

http://openstudy.com/code-of-conduct Read this..

OpenStudy (anonymous):

For that equation that I have written above : General Solution is given by : \[y (I.F) = \int\limits (Q \times I.F ) dt\]

OpenStudy (anonymous):

And I.F is given by : \[\large I.F = e^{\int\limits P. dt}\]

OpenStudy (nirmalnema):

dy/dt=2y dy/2y=dt integrating both side..

OpenStudy (anonymous):

Yes, for sure, you can use here, Variable Separable Method too..

OpenStudy (nirmalnema):

yup...

OpenStudy (anonymous):

you carry on then.. :)

OpenStudy (anonymous):

thank u guys!! :)